# Comparison of Extended Lubrication Theories for Stokes Flow

### v3

@@ Appendix B The Stokes Equations
(36) $U=\frac{\partial\Psi}{\partial Y}\hskip 30.00005ptV=-\frac{\partial\Psi}{\partial X},$
the equations Eqs. 34 and 35 are expressed as,
@@
When $\text{Re}=0$, the formulation Eq. 37 reduces to the Stokes equation $\nabla^{4}\Psi=0$.
To solve the bi-harmonic Navier-Stokes equations, we assume the inlet and outlet flow profiles correspond to a fully developed laminar flow with flux $\mathcal{Q}$. The surface boundary conditions for velocity are Eqs. 6 and 7, and the pressure satisfies Eq. 15. The inlet and outlet velocity profiles are,
@@
The solution $\Psi$, $U$, $V$ to the biharmonic Navier-Stokes equations Eqs. 37 and 36 is determined through an iterative second-order accurate finite difference method as given in [biswas_hoc_2017, gupta_new_2005]. The biharmonic equation Eq. 37 is discretized as,
(43) $28\Psi_{i,j}-8\big(\Psi_{i-1,j}+\Psi_{i+1,j}+\Psi_{i,j-1}+\Psi_{i,j+1}\big)\\ +\big(\Psi_{i-1,j-1}+\Psi_{i-1,j+1}+\Psi_{i+1,j-1}+\Psi_{i+1,j+1}\big)\\ =3\Delta X\big(U_{i,j-1}-U_{i,j+1}+V_{i+1,j}-V_{i-1,j}\big)\\ +\frac{1}{2}\Delta X^{2}\text{Re}\Big(V_{i,j}\big(U_{i-i,j}+U_{i+1,j}+U_{i,j-1}+U_{i,j+1}\big)\\ -U_{i,j}\big(V_{i-1,j}+V_{i+1,j}+V_{i,j-1}+V_{i,j+1}\big)\Big)$
@@
Once the stream and velocity have sufficiently converged, the pressure partial derivatives are determined using a centered second-order accurate finite difference discretisation of the Navier-Stokes equations Eqs. 1 and 2. For Eq. 1,
(46) $\frac{\partial P}{\partial X}\Big|_{i,j}=\frac{1}{\Delta X^{2}}\Big(U_{i-1,j}+U_{i+1,j}-4U_{i,j}+U_{i,j-1}+U_{i,j+1}\Big)\\ -\frac{\text{Re}}{2\Delta X}\Big(U_{i,j}(U_{i+1,j}-U_{i-1,j})+V_{i,j}(U_{i,j+1}-U_{i,j-1})\Big),$
and $\frac{\partial P}{\partial Y}$ as in Eq. 2 is similar. The pressure $P(X,Y)$ is then determined numerically by a path integral from the outlet with the boundary condition $P(1,0)=0$. The average pressure drop $\Delta P_{\text{Stokes}}$ is determined by numerical integration according to Eq. 9 for a given $H$.

### v4

@@ Appendix B The Stokes Equation
(37) $U=\frac{\partial\Psi}{\partial Y},\hskip 30.00005ptV=-\frac{\partial\Psi}{\partial X},$
and assuming a steady state flow, the equations Eqs. 35 and 36 may be expressed as,
@@
The formulation Eqs. 38 and 37 presents an effective method of solution to the Navier-Stokes equations for low to moderate Reynolds numbers [gupta_new_2005, marner_potential-velocity_2014, sen_4oec_2015, biswas_hoc_2017]. Here, we restrict to $\text{Re}=0$, in which case Eq. 38 reduces to the Stokes equation $\nabla^{4}\Psi=0$.
To solve the Stokes equation, we assume the inlet and outlet flow profiles correspond to a fully developed laminar flow with flux $\mathcal{Q}$. The surface boundary conditions for velocity are Eqs. 6 and 7, and the pressure satisfies Eq. 16. The inlet and outlet velocity profiles are,
@@
The solution $\Psi$, $U$, $V$ to the Stokes equation is determined through an iterative second-order accurate finite difference method. The method presented here is adapted from the method of [biswas_hoc_2017, gupta_new_2005] for the full biharmonic Navier-Stokes equations Eqs. 38 and 37. The Stokes equation $\nabla^{4}\Psi=0$ is discretized as,
(44) $28\Psi_{i,j}-8\big(\Psi_{i-1,j}+\Psi_{i+1,j}+\Psi_{i,j-1}+\Psi_{i,j+1}\big)\\ +\big(\Psi_{i-1,j-1}+\Psi_{i-1,j+1}+\Psi_{i+1,j-1}+\Psi_{i+1,j+1}\big)\\ =3\Delta X\big(U_{i,j-1}-U_{i,j+1}+V_{i+1,j}-V_{i-1,j}\big)$
@@
In [dennis_separation_2026], we present a method for approximating the stream and velocity boundary conditions on non-rectilinear domains.
Once the stream and velocity have sufficiently converged, the pressure partial derivatives are determined using a centered second-order accurate finite difference discretisation of the dimensionless Navier-Stokes equations Eqs. 35 and 36 at $\text{Re}=0$. For Eq. 35,
(47) $\frac{\partial P}{\partial X}\Big|_{i,j}=\frac{1}{\Delta X^{2}}\Big(U_{i-1,j}+U_{i+1,j}-4U_{i,j}+U_{i,j-1}+U_{i,j+1}\Big),$
and $\frac{\partial P}{\partial Y}$ as in Eq. 36 is similar. The pressure $P(X,Y)$ is then determined numerically by a path integral from the outlet with the boundary condition $P(1,0)=0$. The pressure drop $\Delta P_{\text{Stokes}}$ is determined by numerical integration according to Eq. 9.